The period of a mass-on-spring oscillator is 3 seconds. Upon decreasing its mass by 500 g, the period will become 2 seconds.

a) Calculate the original mass!

b) Calculate the spring constant! 


Could you please explain in detail how to use the equation below?

"T=2\pi√m/k"


1
Expert's answer
2021-06-14T17:26:43-0400

a)

"T=2\pi \sqrt{\frac{m}{k}},\implies" "k=\frac{4\pi^2m}{T^2},"

"T_1=2\pi \sqrt{\frac{m-m_1}{k}},\implies"

"m=\frac{m_1T^2}{T^2-T_1^2}=\frac{0.5\cdot 9}{9-4}=0.9~kg,"

b)

"k=\frac{4\cdot 3.14^2\cdot 0.9}{9}=3.95~\frac{N}{kg}."